# Twisted crossed products of Banach algebras

### v1

@@ 4. Universal Properties
[…]
@@
Since $A$ has a cai, the nondegeneracy and strong continuity claims follow as in the untwisted case (see for instance Propositions 6.4 and 6.5 in [11]). The fact that the image of $(\lambda^{\mathcal{R}}_{G},\rho_{G}^{\mathcal{R}})$ is in $\mathrm{Iso}(F_{\mathcal{R}}(G,A,\alpha,\sigma))$ follows at once from Lemma 4.6 and the discussion preceding the statement of this Proposition.
@@
Given the equalities (2), (5) from Lemma 4.2, and the equalities (2), (5) from Lemma 4.4, the proof will go exactly as in the untwisted case (see Theorem 2.1 in [9] and Proposition 7.1 in [11] for the details). ∎
@@ 5. Twisted $L^{p}$-operator crossed products
In this section, following [23], we restrict to second countable locally compact groups. This in turn allows us to apply measure theory results such as the Tonelli-Fubini Theorem, Hölder’s inequality, and the Lebesgue Dominated Convergence Theorem to the Lebegue-Bochner spaces $L^{p}(\mu;E)$ from Definition 2.4.
@@
Lemma 2.11 in [24] shows that $\pi$ is indeed a contractive nondegenerate representation of $A$ on $E$. It remains to check Conditions (2), (3), and (4) in Definition 3.8. For Condition (2), we first check that $u_{y}\in\mathrm{Iso}(E)$ for any $y\in G$. Indeed, fix $y\in G$, and using that $\pi_{0}$ is contractive, $\alpha$ isometric, and $\sigma_{x,y}$ has norm 1, we obtain, for any $\xi\in C_{c}(G,L^{p}(\Omega,\mu))\subseteq E$,
@@ 6. Packer-Raeburn trick: $p$-analogue
$\|f^{(\theta)}\|_{\mathcal{R_{\theta}}}=\sup\{\|(\pi\rtimes v)(f^{(\theta)})\|\colon(\pi,v)\in\mathcal{R}_{\theta}\}=\sup\{\|(\pi\rtimes u)(f)\|\colon(\pi,u)\in\mathcal{R}\}=\|f\|_{\mathcal{R}}$
@@
  • AA a separable and nondegenerate $L^{p}$-operator algebra with a cai.
For notational convenience we think of $A$ as a norm closed subalgebra of $\mathcal{B}(L^{p}(\mu))$ with $L^{p}(\mu)$ separable. We also fix the following notation
@@
Let $p\in(1,\infty)$, let $A\subseteq\mathcal{B}(L^{p}(\mu))$ be a separable nondegenerate $L^{p}$-operator algebra with a cai, and let $(G,A,\alpha,\sigma)$ be a twisted dynamical system. Then there is an action $\beta$ of $G$ on $\mathcal{K}_{G}^{p}\otimes_{p}A$ such that $\mathcal{K}_{G}^{p}\otimes_{p}F^{p}(G,A,\alpha,\sigma)$ is isometrically isomorphic to the untwisted crossed product $F^{p}(G,\mathcal{K}_{G}^{p}\otimes_{p}A,\beta)$.

### v2

@@ 4. Universal Properties
In what follows we let $\mathcal{R}$ be a fixed nonempty and uniformly bounded class of covariant representations of a twisted Banach algebra dynamical system $(G,A,\alpha,\sigma)$. Our main goal in this section is to define a “universal” covariant representation of $(G,A,\alpha,\sigma)$ on the Banach space $F_{\mathcal{R}}(G,A,\alpha,\sigma)$. In order to do so, we need the extra assumption that $\mathcal{R}$ consists only of contractive representations. This is only needed to faithfully represent $M(F_{\mathcal{R}}(G,A,\alpha,\sigma))$ on $F_{\mathcal{R}}(G,A,\alpha,\sigma)$, so we will introduce the extra assumption of $C_{\mathcal{R}}\leq 1$ right before is needed.
@@
Since $A$ has a cai, the nondegeneracy and strong continuity claims follow as in the untwisted case (see for instance [11, Propositions 6.4 and 6.5]). The fact that the image of $(\lambda^{\mathcal{R}}_{G},\rho_{G}^{\mathcal{R}})$ is in $\mathrm{Iso}(F_{\mathcal{R}}(G,A,\alpha,\sigma))$ follows at once from Lemma 4.6 and the discussion preceding the statement of this Proposition.
@@
Given the equalities (2), (5) from Lemma 4.2, and the equalities (2), (5) from Lemma 4.4, the proof will go exactly as in the untwisted case (see [9, Theorem 2.1] and [11, Proposition 7.1] for the details). ∎
@@ 5. Twisted $L^{p}$-operator crossed products
In this section, following [23], we restrict to second countable locally compact groups. This in turn allows us to apply measure theory results such as the Tonelli-Fubini Theorem, Hölder’s inequality, and the Lebesgue Dominated Convergence Theorem to the Lebesgue-Bochner spaces $L^{p}(\mu;E)$ from Definition 2.6.
@@
It follows from [24, Lemma 2.11] that $\pi$ is indeed a contractive nondegenerate representation of $A$ on $E$. It remains to check Conditions (2), (3), and (4) in Definition 3.8. For Condition (2), we first check that $u_{y}\in\mathrm{Iso}(E)$ for any $y\in G$. Indeed, fix $y\in G$, and using that $\pi_{0}$ is contractive, $\alpha$ isometric, and $\sigma_{x,y}$ has norm 1, we obtain, for any $\xi\in C_{c}(G,L^{p}(\Omega,\mu))\subseteq E$,
@@ 6. Packer-Raeburn trick: $p$-analogue
$\|f^{(\theta)}\|_{\mathcal{R_{\theta}}}$ $=\sup\{\|(\pi\rtimes v)(f^{(\theta)})\|\colon(\pi,v)\in\mathcal{R}_{\theta}\}$
$=\sup\{\|(\pi\rtimes u)(f)\|\colon(\pi,u)\in\mathcal{R}\}=\|f\|_{\mathcal{R}}$
@@
  • AA a nondegenerate separably representable $L^{p}$-operator algebra with a cai.
For notational convenience (see Notation 2.3) we think of $A$ as a norm closed subalgebra of $\mathcal{B}(L^{p}(\mu))$ with $L^{p}(\mu)$ separable. We also fix the following notation
@@
Let $p\in(1,\infty)$, let $A$ be a nondegenerate separably representable $L^{p}$-operator algebra with a cai, and let $(G,A,\alpha,\sigma)$ be a twisted dynamical system. Then there is an action $\beta$ of $G$ on $\mathcal{K}_{G}^{p}\otimes_{p}A$ such that $\mathcal{K}_{G}^{p}\otimes_{p}F^{p}(G,A,\alpha,\sigma)$ is isometrically isomorphic to the untwisted crossed product $F^{p}(G,\mathcal{K}_{G}^{p}\otimes_{p}A,\beta)$.